OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3

OpenStudy (danjs):

multiply the second equation through by 3, then add the two equations together, see what you get from that

OpenStudy (danjs):

Or from inspection, you can see that the coefficients on the first equation, are 3 times the coefficients in the second, If the coefficients on X and Y are scalar multiples of one another, you have parallel lines

OpenStudy (danjs):

If you put them in y = m*x + b form, you will see they both have the same slope m, but different y intercepts, b, parallel lines.

OpenStudy (danjs):

Solutions are the intersection points (x,y). How many solutions would two parallel lines have then?

OpenStudy (anonymous):

Sooo does that mean there's no solution since they're parallel?

OpenStudy (danjs):

that is right

OpenStudy (danjs):

parallel line - no solution intersecting lines - One Solution overlapping lines - Infinite Solution

OpenStudy (danjs):

Notice if you try and solve these: 3x – 9y = 0 –x + 3y = –3 multiply second by 3 ------------ 3x - 9y = 0 -3x+9y = -9 add the two equations together --------------------- 0x + 0y = -9 0 = -9 ?? FALSE Statement, NO Solutions

OpenStudy (danjs):

sorry reload page

OpenStudy (anonymous):

OOH okay thank you!!